evaluate the determinate
|8 0 0|
|-16 7 8|
|8 4 5|

Answer by scianci(186)

Do 8 times the determinate of
|7 8|
|4 5|
[the matrix that remains after eliminating the row and column to which the 8 belongs in the original matrix]
8(7(5) - 4(8)) =
8(35 - 32) =
8(3) = 24